# Low-altitude Friendly-Jamming for Satellite-Maritime Communications via Generative AI-enabled Deep Reinforcement Learning

### v1

@@ Abstract
Low Earth Orbit (LEO) satellites can be used to assist maritime wireless communications for data transmission across wide-ranging areas. However, extensive coverage of LEO satellites, combined with openness of channels, can cause the communication process to suffer from security risks. This paper presents a low-altitude friendly-jamming LEO satellite-maritime communication system enabled by a unmanned aerial vehicle (UAV) to ensure data security at the physical layer. Since such a system requires trade-off policies that balance the secrecy rate and energy consumption of the UAV to meet evolving scenario demands, we formulate a secure satellite-maritime communication multi-objective optimization problem (SSMCMOP). In order to solve the dynamic and long-term optimization problem, we reformulate it into a Markov decision process. We then propose a transformer-enhanced soft actor critic (TransSAC) algorithm, which is a generative artificial intelligence-enable deep reinforcement learning approach to solve the reformulated problem, so that capturing global dependencies and diversely exploring weights. Simulation results demonstrate that the TransSAC outperforms various baselines, and achieves an optimal secrecy rate while effectively minimizing the energy consumption of the UAV. Moreover, the results find more suitable constraint values for the system.
@@ 1 Introduction
In recent years, with the rapid expansion of the maritime economy, the importance of maritime communications has increased significantly [1]. A reliable and stable communication network is essential to ensure the efficiency of maritime operations [2]. However, the challenges of deploying foundational communication infrastructures, combined with the complexity and variability of the maritime channels, may result in lower transmission rates for maritime networks than terrestrial cellular networks [3]. In this case, satellites, with extensive coverage capabilities, are increasingly being utilized for maritime data transmission, such that facilitating the effective exchange of data from vessels at sea [4]. In particular, the low Earth Orbit (LEO) satellites, operating closer to the Earth, have higher stability for enhancing communication performance [5][6]. However, the open channels and extensive coverage of LEO satellites make them vulnerable to unauthorized access and potential eavesdropping by malicious users, which may pose security risks [7]. Although the traditional cryptography-based methods can cope with the security risks in some cases, they require frequent data encoding and decoding, and complex key distribution and management, which pose challenges in the energy-limited maritime environment when large amounts of data are transmitted [8].
@@
In this case, physical layer security (PLS) can dynamically adjust security mechanisms, making it an effective way to ensure secure communications in the LEO satellite-maritime networks [9]. For example, through intelligent beamforming methods, LEO satellites can focus signals on target vessels by optimizing the directionality and power distribution of the transmitting antenna, thereby reducing the probability of illegitimate users acquiring signals [10]. However, for high-speed moving satellites, real-time computation and adjustment of beamforming parameters place high demands on the hardware, leading to a greater computational burden [11]. As such, low-altitude platforms are used to introduce friendly-jamming signals to enhance PLS.
@@
Common low-altitude platforms include unmanned aerial vehicles (UAVs) and electric vertical take-off and landing (eVTOL) aircraft [12]. In complex maritime environments, UAVs, with adjustable flight altitudes and 3D mobility, are particularly effective at overcoming movement limitations imposed by sea terrain and obstacles [13]. For example, the authors in [14] utilized UAVs to form a virtual antenna array as a jammer to send jamming signals to the illegitimate vessel, thus protecting legitimate data signals against being eavesdropped. The authors in [15] introduced a cooperative jamming scheme for UAV-assisted networks to elevate security by regulating the position of UAVs. However, the aforementioned works overlook the mobility of vessels, which is a crucial factor for maintaining reliable connectivity in practical scenarios [16]. In such dynamic conditions, the original solutions are no longer effective, and also calculating solutions in real-time while in operation can be time-consuming. Thus, it is essential to account for the dynamics of realistic scenarios when implementing LEO satellite-maritime PLS via UAVs.
@@
Implementing such systems in the considered scenario encounters several challenges. First, the movements of vessels are highly dynamic as they traverse the ocean in varying directions, constantly changing the relative positions in the communication networks. Similarly, LEO satellites experience rapid motion, continuously altering their positions relative to the vessels. In this case, traditional offline methods (e.g., convex optimization and evolutionary computation) are inadequate for addressing the problem in such a dynamic condition [17][18]. Second, when controlling UAVs for jamming, the potential impact on legitimate users needs to be considered to ensure their normal communications. As such, the transmit power of the UAV should be carefully optimized based on the requirements of different users. Finally, during the optimization, the UAVs need to frequently regulate the positions according to vessel movements, which increases energy consumption. Thus, secure communication performance and energy efficiency are conflicting objectives and difficult to be balanced. Moreover, the relative importance of these objectives can change depending on the specific scenarios, rendering existing methods (e.g., [19][20]) that prioritize one goal and constrain others inapplicable for this case. Therefore, an innovative approach, different from previous work, is needed to optimize multiple conflicting objectives in dynamic environments simultaneously.
@@
Accordingly, we formulate a multi-objective optimization problem (MOP), and propose a generative AI (GenAI)-enabled online approach to solve the problem. The main contributions of this work are summarized as follows.
@@
  • Low-altitude Friendly-Jamming LEO Satellite-Maritime Communication System: We present a low-altitude friendly-jamming LEO satellite-maritime communication system enabled by a UAV, where the LEO satellite sends data signals to the legitimate vessel within range, and the UAV sends jamming signals to the potential eavesdropping vessel. This system is the first to holistically consider movements of LEO satellites and vessels, and design a dynamic low-altitude friendly-jamming mechanism to ensure secure real-time data transmission.
@@
  • Multi-objective Optimization Problem Formulation: The security of LEO satellite-maritime communications and energy efficiency of the system are in conflict. Accordingly, we formulate a secure satellite-maritime communication multi-objective optimization problem (SSMCMOP) to simultaneously maximize the secrecy rate and minimize the energy consumption of the UAV. However, the problem is an NP-hard and long-term optimization problem, which makes it more complicated to be solved.
@@
  • GenAI-enabled Deep Reinforcement Learning (DRL) Approach: The conventional DRL algorithms confront the challenges of large solution spaces and complex nonlinear relationships when addressing the SSMCMOP. In this case, we first reformulate the problem into a Markov decision process (MDP). We then propose a transformer-enhanced soft actor critic (TransSAC) algorithm, which is a GenAI-enabled DRL approach to solve the problem. Specifically, TransSAC enhances the learning ability of the long-term problem by capturing global dependencies, and explores diverse weights to effectively balance multiple optimization objectives.
@@
  • Simulations and Performance: Simulation results indicate that the proposed UAV-assisted approach can accomplish secure and energy-efficient LEO satellite-maritime communications, and outperforms the non-UAV approach. Moreover, comparative results illustrate that the TransSAC outperforms other conventional DRL algorithms, achieving an optimal secrecy rate with minimum energy consumption of the UAV. In addition, we identify optimal constraint values for the formulated MDP, thereby enhancing the performance of algorithms.
@@
The rest of this paper is organized as follows: Section 2 reviews the related work. Section 3 introduces the models and preliminaries. Section 4 formulates the SSMCMOP. The GenAI-enabled DRL approach is proposed in Section 5. Section 6 illustrates the simulation results, and Section 7 summarizes the overall work.
@@ 2.1 LEO Satellite-Maritime Communications
The rapid advancement of maritime wireless communications has attracted much attention, while deployment problems in practice pose a considerable challenge to transmission rates [21]. The satellites, with wide coverage capabilities, serve as valuable auxiliary equipment to facilitate the effective exchange of centralized data and information between vessels at sea [22]. For example, the authors in [23] considered establishing a service-driven integrated satellite-terrestrial maritime communication network (MCN) to improve transmission efficiency, extend network range, and provide maritime-specific services. Moreover, the authors in [24] proposed an intelligent spectrum-sharing scheme for the satellite-maritime integrated network to optimize throughput and spectral efficiency. However, the significant latency caused by long-distance satellite transmission can negatively impact the efficiency of communications.
LEO satellites, with their proximity to Earth, significantly reduce latency and improve data transmission efficiency. Moreover, LEO satellites can be deployed in satellite constellations to achieve seamless global coverage, allowing them to be gradually applied in maritime communications. For example, the authors in [25] proposed an LEO satellite-assisted shore-to-vessel network to achieve end-to-end communications. In addition, the authors in [26] introduced a hybrid LEO and UAV edge computing method in space–air–sea integrated networks for marine Internet of Things (IoT) systems, and demonstrated the energy savings. Note that the open channels and extensive coverage of satellites make them susceptible to eavesdropping during data transmission. However, the aforementioned works focus mainly on communication efficiency, which ignores the potential security risks.
@@ 2.2 Security Strategies
As aforementioned, it is essential to take measures to improve the security of maritime communications [27]. As such, the authors in [28] presented analyses on cross-layer attacks and security measures in the satellite MCNs. In addition, the authors in [29] explored a secure and robust communication scheme by modeling the channel phase and angle uncertainties of geostationary Earth orbit and LEO. However, the encryption and decryption methods in the abovementioned works have limitations, as transmitting large amounts of data requires significant computational energy, resulting in transmission delays.
In this case, PLS can dynamically adapt security mechanisms according to the channel conditions, achieving secure transmissions for the maritime networks [13]. For example, the authors in [30] aimed to enhance the performance of the satellite-terrestrial MCN by optimizing the transmission beamforming of the base station and LEO satellites. However, for fast-moving LEO satellites, the immediate processing and regulating of beamforming parameters demand significant resources, increasing computational load. As such, low-altitude platforms are employed to introduce friendly jamming signals, which can significantly enhance PLS [14]. Common low-altitude platforms, such as an eVTOL aircraft and UAVs, play key roles in various applications [12]. Moreover, UAVs, with high mobility and flexible deployment, are well-suited for implementing maritime PLS [31]. For example, the authors in [32] considered a power allocation scheme to optimize secrecy throughput to improve the PLS of downlink transmission under the UAV energy constraint. In addition, the authors in [33] presented a reinforcement learning-enabled UAV maritime communication relay strategy via a deep neural network with a dueling structure to resist jamming attacks. However, the aforementioned works overlook the complex trade-offs between conflicting objectives such as safety performance and energy efficiency, as well as the varying importance of these objectives across different scenarios, which often leads to suboptimal solutions.
@@ 2.3 Optimization Approaches
To address these challenges, MOP provides a mathematical framework for optimizing multiple conflicting objectives simultaneously [34]. By formulating an MOP, we can model the correlations between conflicting objectives and find solutions that offer a better compromise across varying conditions. Generally, several common methods exist for handling MOP. First, swarm intelligence optimization algorithms can solve MOP, which gradually approximates the Pareto front and obtains multiple non-dominated solutions [35]. For example, the authors in [36] proposed a collaborative beamforming method to resist illegitimate eavesdropping via UAVs, and presented an improved swarm intelligence algorithm to address MOP challenges. Moreover, the authors in [37] considered a UAV-assisted remote communication scenario while multiple eavesdroppers aim to intercept the data, and proposed a multi-objective salp swarm algorithm to deal with the MOP. However, swarm intelligence algorithms lack dynamic adjustment and timely feedback mechanisms, which affects the overall performance, thereby not applicable to real-time problems.
@@
Furthermore, DRL has adaptive learning capabilities and can handle complex action spaces, making it suitable for solving dynamic timing problems [38]. For example, the authors in [39] considered a dual-UAV secure maritime communication system that maximizes the minimum secrecy rate for the mobile user over all time slots. Moreover, the authors in [40] considered a UAV reconfigurable intelligent surface-assisted maritime communication system to maximize energy efficiency and ensure the quality of requirements against jamming attacks. However, the aforementioned works ignore the possible local optimal problem in long-term optimization, as well as generate suboptimal solutions due to improper weight value settings.
In summary, different from previous works, we employ low-altitude friendly-jamming LEO satellite-maritime communications enabled by a UAV and propose a novel GenAI-enable DRL approach that can derive a high-quality policy for resolving dynamic SSMCMOP.
@@ 3 Models and preliminaries
In this section, we first present the low-altitude friendly-jamming LEO satellite-maritime communication system. Then, we introduce the LEO satellite orbit model and vessel movement model. Next, the corresponding communication network is presented. Finally, the energy consumption model of the UAV is presented. In addition, the main notations are shown in Table I.
@@ 3.1 System Overview
As shown in Fig. 1, we consider a low-altitude friendly-jamming LEO satellite-maritime communication system, involving an LEO satellite, a UAV, a legitimate vessel denoted as Alice, and an illegitimate user denoted as Eve. Among them, the LEO satellite receives information from the data fusion center and forwards it to the target vessel Bob within range. Note that the LEO satellite is equipped with high-performance antennas with adequate power, which can ensure efficient downlink communication to the vessels. However, its open channel is vulnerable to eavesdropping by Eve. In this case, due to the dynamics of the marine channels and constraints of marine routes, it is challenging to combat the eavesdropping attack with existing vessels. Moreover, the offshore evaporation ducts and multipath scattering increase the complexity of vessel-aided methods. Therefore, we mainly focus on the low-altitude friendly-jamming platforms to deal with eavesdropping attack. Among the platforms, UAVs, with great mobility, high flexibility, and wide coverage, are particularly suited for implementing secure maritime communications [41].
Figure 1: A low-altitude friendly-jamming LEO satellite-maritime communication system.
@@
During the process of satellite-maritime communications, we utilize the 3D Cartesian coordinate system to indicate the time-varying positions of Alice, Eve, the UAV, and the LEO satellite at time slot $t$ as ($x_{A}[t]$, $y_{A}[t]$, $z_{A}[t]$), ($x_{E}[t]$, $y_{E}[t]$, $z_{E}[t]$), ($x_{U}[t]$, $y_{U}[t]$, $z_{U}[t]$), and ($x_{S}[t]$, $y_{S}[t]$, $z_{S}[t]$), respectively. The satellite spectrum resources are limited and UAV needs to share spectrum with satellites, which complicates the dynamic communications [42]. To this end, we model the LEO satellite orbit and the vessel movement, as well as the communication processes of the LEO satellite and vessels, the UAV and vessels, to express the dynamics of the system.

### v2

# Low-altitude UAV Friendly-Jamming for Satellite-Maritime Communications via Generative AI-enabled Deep Reinforcement Learning

@@ Abstract
Low Earth orbit (LEO) satellites can be used to assist maritime wireless communications for wide-area data transmission. However, the extensive coverage of LEO satellites, combined with the openness of channels, can cause the communication process to suffer from security risks. This paper presents a LEO satellite-maritime communication system assisted by low-altitude unmanned aerial vehicle (UAV) friendly-jamming to ensure data security at the physical layer. Since such a system requires balancing the conflicting performance metrics of secrecy rate and energy consumption of the UAV to meet evolving scenario demands, we formulate a secure satellite-maritime communication multi-objective optimization problem (SSMCMOP). In order to solve the dynamic and long-term optimization problem, we reformulate it into a Markov decision process. We then propose a transformer-enhanced soft actor-critic (TransSAC) algorithm, which is a generative artificial intelligence-enabled deep reinforcement learning approach to solve the reformulated problem, thus capturing strong temporal correlations and diversely exploring weights. Simulation results demonstrate that the TransSAC algorithm outperforms comparative approaches and algorithms, maximizing the secrecy rate while effectively minimizing the energy consumption of the UAV. Moreover, the results identify more suitable constraints for the system.
@@ 1 Introduction
In recent years, with the rapid expansion of the maritime economy, the importance of maritime communications has increased significantly [Yuan2024]. A reliable and stable communication network is essential to ensure the efficiency of maritime operations [Zeng2025]. However, the challenges of deploying foundational communication infrastructures, combined with the complexity and variability of the maritime channels, may result in lower transmission rates for maritime networks than terrestrial cellular networks [Zhang2024]. In this case, satellites, with extensive coverage capabilities, are increasingly being utilized for maritime data transmission, thereby facilitating the effective exchange of data from vessels at sea [Zhou2023]. In particular, the low Earth orbit (LEO) satellites, operating closer to the Earth, have higher stability for enhancing communication performance [Pokhrel2024]. However, the open channels and extensive coverage of LEO satellites make them vulnerable to unauthorized access and potential eavesdropping by malicious users, which may pose security risks [Yue2023]. Although the traditional cryptography-based methods can cope with the security risks in some cases, they require frequent data encoding and decoding, and complex key distribution and management, which pose challenges in the energy-limited maritime environment when large amounts of data are transmitted [Wu2018].
@@
In this case, physical layer security (PLS) can dynamically adjust security mechanisms, making it a suitable way to ensure secure communication in the LEO satellite-maritime networks [Wang2019]. For example, through intelligent beamforming methods, LEO satellites can focus signals on target vessels by optimizing the directionality and power distribution of the transmitting antenna, thereby reducing the probability of illegitimate users acquiring signals [Zheng2022]. However, for high-speed moving LEO satellites, real-time computation and adjustment of beamforming parameters place high demands on the hardware, leading to a greater computational burden [Xv2024]. As such, low-altitude platforms are used to introduce friendly-jamming signals to enhance PLS.
@@
Common low-altitude platforms include unmanned aerial vehicles (UAVs) and electric vertical take-off and landing (eVTOL) aircraft [Huang2024b]. In complex maritime environments, UAVs, with low costs, adjustable flight altitudes, and flexible mobility, are particularly suitable for overcoming movement and operation limitations imposed by sea terrain and obstacles [Qian2022]. For example, the authors in [Huang2025] utilized UAVs to form a virtual antenna array as a jammer to send jamming signals to an illegitimate vessel, thus protecting legitimate data signals against being eavesdropped. The authors in [DangNgoc2022] introduced a cooperative jamming scheme for UAV-assisted networks to enhance security by regulating the position of UAVs. However, the aforementioned works overlook the mobility of vessels, which is a crucial factor for maintaining reliable connectivity in practical scenarios [Wang2024]. In such dynamic conditions, the original solutions become ineffective, as vessel movement degrades previously stable links. Moreover, continuous real-time computation during the operation creates computational overhead that affects the responsiveness of the system. Therefore, to achieve the LEO satellite-maritime PLS communication system via UAVs, we need to focus on the movement of vessels and the temporal dynamics of maritime channels.
@@
Implementing such systems encounters several challenges. First, the movements of vessels are dynamic as they traverse the ocean in varying directions, constantly changing their relative positions in communication networks. Similarly, LEO satellites experience rapid motion, continuously altering their positions relative to the vessels. In this case, traditional offline methods (e.g., convex optimization and evolutionary computation) become ineffective in such a dynamic condition [Yuan2020][Wang2023], while deep reinforcement learning (DRL) algorithms integrate the feature extraction capabilities of deep learning and the decision-making capabilities of reinforcement learning, demonstrating remarkable adaptability in dynamic environments [Liao2024][Shao2024]. Second, when controlling UAVs for jamming, the potential impact on legitimate users needs to be considered to ensure their normal communications. As such, the transmit power of the UAV should be carefully optimized based on user requirements. Meanwhile, UAVs need to frequently adjust their positions to optimize system performance, which increases their energy consumption. Thus, secure communications and energy consumption are conflicting and require balancing. Additionally, since the relative importance of these objectives can change depending on the specific scenarios, existing methods (i.e.[Cai2020]) that prioritize one goal and constrain others become inapplicable for this case. Finally, our considered scenario is time-sensitive and involves large-scale decision making, which requires solutions that can adapt quickly while remaining computationally efficient, further increasing the complexity. Thus, an innovative approach, different from previous works, is needed to jointly optimize multiple conflicting objectives and efficiently respond to the dynamic maritime conditions.
@@
Accordingly, we formulate a multi-objective optimization problem (MOP), and propose a generative AI (GenAI)-enabled DRL approach to solve the problem. The main contributions of this work are summarized as follows.
@@
  • Low-altitude UAV Friendly-Jamming for the LEO Satellite-Maritime Communication System: We present an LEO satellite-maritime communication system assisted by low-altitude UAV friendly-jamming, where an LEO satellite sends data signals to a legitimate vessel within range, and a UAV sends jamming signals to a potential eavesdropping vessel. This system is the first to holistically consider the movements of LEO satellites and vessels while designing an adaptive UAV friendly-jamming mechanism that ensures real-time maritime communication security.
@@
  • Multi-objective Optimization Problem Formulation: The security performance and energy consumption of the system are in conflict due to shared decision variables. To balance these key performance metrics, we formulate a secure satellite-maritime communication multi-objective optimization problem (SSMCMOP) to jointly maximize the secrecy rate and minimize the energy consumption of the UAV. However, the problem is an NP-hard and long-term optimization problem, making it more complex to be solved.
@@
  • GenAI-enabled DRL Approach: Conventional DRL algorithms confront the challenges of large solution spaces and strong temporal correlation when addressing the formulated SSMCMOP. In this case, we first reformulate the problem into a Markov decision process (MDP). We then propose a transformer-enhanced soft actor-critic (TransSAC) algorithm, which is a GenAI-enabled DRL approach to solve the problem. Specifically, the TransSAC algorithm captures temporal dependencies through encoded sequences and accelerates learning policies via parallel processing, while diversely exploring weights to effectively balance multiple optimization objectives.
@@
  • Simulations and Performance: Simulation results indicate that the proposed UAV-assisted approach can achieve secure and energy-efficient LEO satellite-maritime communications, significantly outperforming the non-UAV approach. Moreover, comparative results illustrate that TransSAC outperforms other conventional DRL algorithms, achieving the maximum secrecy rate with the minimum energy consumption of the UAV. In addition, we identify optimal constraint values for the formulated MDP, thereby enhancing the performance of algorithms.
@@
The rest of this paper is organized as follows: Section 2 reviews the related work. Section 3 introduces the models and preliminaries. Section 4 formulates the SSMCMOP. The GenAI-enabled DRL approach is proposed in Section 5. Section 6 illustrates the simulation results. Section 7 presents some discussions and Section 8 summarizes the overall work.
@@ 2.1 LEO Satellite-Maritime Communications
The rapid advancement of maritime wireless communications has attracted much attention, while deployment problems in practice pose a considerable challenge to transmission rates [He2025]. The satellites, with wide coverage capabilities, serve as valuable auxiliary equipment to facilitate the effective exchange of centralized data and information between vessels at sea [Alqurashi2023]. For example, the authors in [Dai2025] proposed an energy-efficient multi-access edge computing scheme for heterogeneous satellite-maritime networks to enhance the perception and offloading endurance of UAVs. Moreover, the authors in [Wu2024] proposed an intelligent spectrum-sharing scheme for the satellite-maritime integrated network to optimize throughput and spectral efficiency. However, the significant latency caused by long-distance satellite transmission can negatively impact the efficiency of communications.
LEO satellites, with their proximity to Earth, significantly reduce latency and improve data transmission efficiency. Moreover, LEO satellites can be deployed in satellite constellations to achieve seamless global coverage, allowing them to be gradually applied in maritime communications. For example, the authors in [Hu2024] proposed an LEO satellite-assisted shore-to-vessel network to achieve end-to-end communications. The authors in [Senadhira2025] considered a UAV-assisted LEO satellite communication network to provide coverage for low-end maritime users. Note that the open channels and extensive coverage of satellites make them susceptible to eavesdropping during data transmission. However, the aforementioned works focus mainly on communication efficiency, which ignores the potential security risks.
@@ 2.2 Security Strategies
As aforementioned, it is essential to take measures to improve the security of maritime communications. As such, the authors in [Guo2022] presented analyses on cross-layer attacks and security measures in the satellite MCNs. In addition, the authors in [Jiang2023] explored a secure and robust communication scheme by modeling the channel phase and angle uncertainties of geostationary Earth orbit and LEO. However, the encryption and decryption methods in the abovementioned works have limitations, as transmitting large amounts of data requires significant computational energy, resulting in transmission delays.
In this case, PLS can dynamically adapt security mechanisms according to the channel conditions, achieving secure transmissions for the maritime networks [Qian2022]. For example, the authors in [Xiong2024] aimed to enhance the performance of the satellite-terrestrial MCN by optimizing the transmission beamforming of the base station and LEO satellites. However, for fast-moving LEO satellites, the immediate processing and regulating of beamforming parameters demand significant resources, increasing computational load. As such, low-altitude platforms are employed to introduce friendly jamming signals, which can significantly enhance PLS [Huang2025]. Common low-altitude platforms, such as eVTOL aircrafts and UAVs, play key roles in various applications [Huang2024b]. Moreover, UAVs, with low cost, high mobility, and flexible deployment, are well-suited for implementing maritime PLS [Qian2022]. For example, the authors in [Wang2019a] considered a power allocation scheme to optimize secrecy throughput to improve the PLS of downlink transmission. In addition, the authors in [Liu2022] presented a reinforcement learning-enabled UAV maritime communication relay strategy with a dueling structure to resist jamming attacks. However, the aforementioned works overlook the complex trade-offs between conflicting objectives such as security performance and energy efficiency, as well as the varying importance of these objectives across different scenarios, which often leads to suboptimal solutions.
@@ 2.3 Optimization Approaches
To address these challenges, MOP provides a mathematical framework for optimizing multiple conflicting objectives simultaneously [Karami2022]. By formulating an MOP, we can model the correlations between conflicting objectives and find solutions that offer a better compromise across varying conditions. Generally, several common methods exist for handling MOP. First, swarm intelligence optimization algorithms can solve MOP, which gradually approximates the Pareto front and obtains multiple non-dominated solutions [Zheng2024]. For example, the authors in [Huang2025] proposed a collaborative beamforming method to resist eavesdropping via UAVs, and presented a swarm intelligence algorithm to address MOP. Moreover, the authors in [Sun2023] considered a UAV-assisted communication scenario where eavesdroppers aim to intercept the data, and proposed a multi-objective salp swarm algorithm to deal with the MOP. However, swarm intelligence algorithms lack dynamic adjustment and timely feedback mechanisms, which affects overall performance and is not applicable to real-time problems.
@@
Furthermore, DRL has adaptive learning capabilities and can handle complex action spaces, making it suitable for solving dynamic timing problems [Li2024]. For example, the authors in [Wang2021] considered a dual-UAV secure maritime communication system that maximizes the minimum secrecy rate for the mobile user over all time slots. Moreover, the authors in [Yang2024] considered a UAV reconfigurable intelligent surface-assisted maritime communication system to maximize energy efficiency and ensure the quality of requirements against jamming attacks. However, the aforementioned works overlook crucial challenges such as the local optimal problem in the long-term optimization process, computational burden in large-scale action spaces, and suboptimal solutions due to improper weight value settings.
In summary, different from previous works, we employ low-altitude UAV friendly-jamming for LEO satellite-maritime communications and propose a novel GenAI-enabled DRL approach that can derive a high-quality policy for resolving dynamic SSMCMOP. This approach is more suitable for dynamic scenarios requiring trade-offs between conflicting objectives while enabling fast response to changing conditions.
@@ 3 Models and preliminaries
In this section, we first present the LEO satellite-maritime communication system assisted by a low-altitude UAV. Then, we introduce the LEO satellite orbit model and vessel movement model. Next, the corresponding communication network is presented. Finally, the energy consumption model of the UAV is presented. In addition, the main notations are shown in Table I.
@@ 3.1 System Overview
As shown in Fig. 1, we consider an LEO satellite-maritime communication system assisted by a low-altitude UAV, which involves an LEO satellite, a UAV, a legitimate vessel denoted as Alice, and an illegitimate user denoted as Eve. Among them, the LEO satellite receives information from the data fusion center and forwards it to the target vessel Bob within range. Note that the LEO satellite is equipped with high-performance antennas with adequate power, which can ensure efficient downlink communication to the vessels. However, its open channel is vulnerable to eavesdropping by Eve. In this case, due to the dynamics of the marine channels and constraints of marine routes, it is challenging to combat the eavesdropping attack with existing vessels. Moreover, the offshore evaporation ducts and multipath scattering increase the complexity of vessel-aided methods. Therefore, we mainly focus on the low-altitude friendly-jamming platforms to deal with eavesdropping attacks. Among the platforms, UAVs, with great mobility, high flexibility, and wide coverage, are particularly suited for implementing secure maritime communications [Liu2022a].
Figure 1: An LEO satellite-maritime communication system assisted by low-altitude UAV.
@@
During the process of satellite-maritime communications, we utilize the 3D Cartesian coordinate system to indicate the time-varying positions of Alice, Eve, the UAV, and the LEO satellite at time slot $t$ as ($x_{A}[t]$, $y_{A}[t]$, $z_{A}[t]$), ($x_{E}[t]$, $y_{E}[t]$, $z_{E}[t]$), ($x_{U}[t]$, $y_{U}[t]$, $z_{U}[t]$), and ($x_{S}[t]$, $y_{S}[t]$, $z_{S}[t]$), respectively. The satellite spectrum resources are limited and UAV needs to share spectrum with satellites, which complicates the dynamic communications [Li2020]. To this end, we model the LEO satellite orbit and the vessel movement, as well as the communication processes of the LEO satellite and vessels, the UAV and vessels, to express the dynamics of the system.